# Towards a Unified View of Large Language Model Post-Training

### v1

@@ 1 Introduction
To circumvent these challenges, recent works have focused on integrating SFT or SFT-style imitation learning losses directly with RL objectives (Yan et al., 2025; Fu et al., 2025; Zhang et al., 2025a). In these approaches, the model is updated using a composite loss function. The balance between the imitation and exploration components is governed by various strategies, including a fixed coefficient, a predefined schedule, a dynamic adjustment based on entropy, or a learnable parameter. These works predominantly treat the SFT and RL losses as two distinct objectives. And a detailed analysis of why these two learning signals can be effectively combined within a unified optimization process remains largely unexplored.
@@
Takeaways 1. UPGE provides a theoretical unification of a wide spectrum of post-training algorithms, covering both SFT and RL losses within a single formulation ($\S$ 3). 2. HPT is capable of outperforming previous post-training and mixed-policy algorithms across a diverse range of models ($\S$ 4). 3. Dynamic integration of SFT and RL in HPT achieves the highest Pass@1024, facilitating enhanced exploration and generalization of the model ($\S$ 5.1).
2 Related Works
2.1 LLM Post-Training: SFT and RL
Current post-training methodologies for LLMs are largely centered around two primary paradigms: Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL) (Wei et al., 2021; Ouyang et al., 2022). In the SFT paradigm, models are adapted for specific applications through training on curated input-output pairs, a process which has been shown to effectively align their behavior with human demonstrations (Chung et al., 2022; Longpre et al., 2023; Touvron et al., 2023a; b). In parallel, numerous works have highlighted RL as an effective approach for refining LLM behavior in ways that are difficult to capture with SFT’s static datasets (Glaese et al., 2022; Bai et al., 2022; Nakano et al., 2021). Within this domain, a popular framework is Reinforcement Learning from Human Feedback (RLHF), which optimizes the LLM policy against a reward model trained on human preferences (Christiano et al., 2017; Stiennon et al., 2020). Multiple works have established Proximal Policy Optimization (PPO) as a cornerstone algorithm for this phase (Schulman et al., 2017; Ziegler et al., 2019). To further improve reasoning capabilities in reward-driven optimization, recent advancements like Group Relative Policy Optimization (GRPO) have also been developed and widely adopted (Shao et al., 2024; Zheng et al., 2025; Chen et al., 2025).
2.2 A Combination of Online and Offline Data in LLM Post-Training
Beyond applying SFT or RL in isolation, further explorations have sought to synergize their respective strengths by combining signals from pre-existing offline data and dynamically generated online data (Fu et al., 2025; Yan et al., 2025). This motivation stems from the distinct characteristics of each approach: SFT is noted for its efficiency in distilling knowledge from offline sources, whereas RL is valued for fostering exploration through online rollouts, a process frequently linked to improved generalization (Rajani et al., 2025; Chu et al., 2025). The strategies for this integration are diverse; some techniques use offline data as a prefix to guide online generation (Zhou et al., 2023; Touvron et al., 2024; Li et al., 2025; Wang et al., 2025), while others enhance offline data by incorporating reward signals in a process known as reward-augmented fine-tuning (Liu et al., 2024; Zhao et al., 2023; Park et al., 2025; Sun et al., 2025). The broader landscape also includes various purely offline preference optimization methods, though they follow a different paradigm (Rafailov et al., 2023; Mitchell et al., 2024; Liu et al., 2025c; Ethayarajh et al., 2024; Ahmadian et al., 2024). However, the most direct approach to synergy involves the concurrent use of both data types for training updates.
This direct approach, often termed mix-policy learning, is particularly relevant to our work and typically involves updating the model with a composite objective that combines an SFT loss from offline data and an RL loss from online data (Dong et al., 2023; Gulcehre et al., 2023; Singh et al., 2023; Liu et al., 2023). For instance, LUFFY (Yan et al., 2025) explores this paradigm by combining a fixed ratio of offline demonstration data with online rollouts in each training batch. Subsequently, SRFT (Fu et al., 2025) proposed a monolithic training phase that dynamically adjusts the weights of SFT and RL losses based on the model’s policy entropy, further demonstrating the viability of unifying these signals over a sequential pipeline. The principle of creating such a composite loss is shared by a variety of other recent frameworks (Wu et al., 2024; Zhang et al., 2025a; Kim et al., 2025; Yu et al., 2025; Liu et al., 2025a), and AMFT (He et al., 2025) begins to explore meta-gradient-based controllers. While these methods highlight a clear trend towards unifying training signals, a foundational theoretical analysis explaining why these different learning signals can be effectively combined is still lacking. This motivates our work to establish a unified theoretical framework that in turn inspires a more principled algorithm design.
@@ 3 A Unified View on Post-Training Algorithms
GRPO
(Shao et al., 2024)
$\pi_{ref}=\pi_{\theta_{old}}$ $\hat{A}_{GRPO}=\frac{R(\tau_{j})-\text{mean}(\{R(\tau_{j})\}_{G_{on}})}{\text{std}(\{R(\tau_{j})\}_{G_{on}})}$ $\nabla\mathcal{J}_{GRPO}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{GRPO}\mathbb{1}_{\text{Clip}}}{\pi_{ref}(\tau)}$
REINFORCE
(Ahmadian et al., 2024)
$\pi_{ref}=\pi_{\theta}$ $\hat{A}_{REINFORCE}=\pm 1$ $\nabla\mathcal{J}_{REF.}(\theta)=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{REF.}}{\pi_{\theta}(\tau)}$
CISPO
(Chen et al., 2025)
$\pi_{ref}=\pi_{\theta_{old}}$ $\hat{A}_{CISPO}=\hat{A}_{GRPO}$ $\nabla\mathcal{J}_{CISPO}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{CISPO}\mathbb{1}_{\text{CIS-Mask}}}{\pi_{ref}(\tau)}$
GSPO
(Zheng et al., 2025)
$\pi_{ref}=\pi_{\theta}\left(\frac{\pi_{\theta_{old}}(\tau_{i,j}|q_{i})}{\pi_{\theta}(\tau_{i,j}|q_{i})}\right)^{1/|\tau_{i,j}|}$ $\hat{A}_{GSPO}=\hat{A}_{GRPO}$ $\nabla\mathcal{J}_{GSPO}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{GSPO}\mathbb{1}_{\text{Seq-Clip}}}{\pi_{ref}(\tau)}$
@@
SRFT (Offline)
(Fu et al., 2025)
$\pi_{ref}\equiv 1$ $\hat{A}_{SRFT}=\frac{R(\tau_{j})-\text{mean}(\{R(\tau_{j})\}_{G_{on}\cup G_{off}})}{\text{std}(\{R(\tau_{j})\}_{G_{on}\cup G_{off}})}$ $\nabla\mathcal{J}_{SRFT}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{SRFT}}{\pi_{ref}(\tau)=1}$
LUFFY (Offline)
(Yan et al., 2025)
$\pi_{ref}\equiv 1$ $\hat{A}_{LUFFY}=\hat{A}_{SRFT}$ $\nabla\mathcal{J}_{LUFFY}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{LUFFY}}{\pi_{ref}(\tau)=1}f_{\text{shape}}^{\prime}$
@@ 3.2 Derivation of the Unified Policy Gradient Estimator
$\widehat{A}_{uni}(\tau,q)=\underbrace{r(\tau\mid q)}_{\widehat{A}_{\mathrm{RL}}(\tau,q)}\;+\;\underbrace{\mu\,\mathbb{1}\{\pi_{ref}=\pi_{\beta}\}\,\frac{\pi_{\beta}(\tau\mid q)}{\pi_{\theta}(\tau\mid q)}}_{\widehat{A}_{\mathrm{SFT}}(\tau,q)}.$ (4)
@@ 3.3 Gradient Component Analysis
Apart from the strong connection to data source and sampling polices, some studies employ a hand-crafted reweight factor within the reference policy denominator. These works (Yan et al., 2025; Zhang et al., 2025b) typically find desirable token properties and purposefully place a higher/lower weight on these desirable/undesirable tokens, respectively.
@@
There are two commonly used settings for estimating the sequence-level advantage function: the fixed advantage setting and the adaptive advantage setting. The fixed setting considers $\hat{A}=\pm 1$ given the rule-based verification, which is adapted by REINFORCE and implicitly by SFT (where all sequences are positive samples). Alternatively, recent studies have focused on using adaptive advantage estimations, performing re-centering or normalization based on the performance of the current rollout group. Notably, GRPO and its variants, such as DAPO (Yu et al., 2025) and LUFFY (Yan et al., 2025), use unit normalization such that the advantage estimation of the group has a unit standard deviation. Other approaches, such as Dr. GRPO (Liu et al., 2025b), RLOO (Ahmadian et al., 2024), and REINFORCE++ (Hu et al., 2025a), claim that dividing the standard deviation introduces a difficulty bias and that only recentering is adequate.
@@ 3.4 Hybrid Post-Training with Performance Feedback
 $P\leftarrow\frac{1}{n}\sum_{i=1}^{n}v(\tau_{i})$$\alpha\leftarrow f(P),\quad\beta\leftarrow g(P)$Performance feedback on question $q$ Compute on-policy RL loss $\mathcal{L}_{\mathrm{RL}}$ using rollouts $\{\tau_{i}\}$ and normalized advantages derived from $\{R(\tau_{i})\}$. Compute SFT loss $\mathcal{L}_{\mathrm{SFT}}$ on the supervising trajectory $\tau^{\star}$. $\mathcal{L}\leftarrow\alpha\,\mathcal{L}_{\mathrm{RL}}+\beta\,\mathcal{L}_{\mathrm{SFT}}$Mixed loss with performance feedback coefficients $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}$
@@ 4.1 Experimental Setup
We evaluate HPT on $6$ mathematical reasoning benchmarks: AIME 2024 (Li et al., 2024), AIME 2025 (Li et al., 2024), AMC (Li et al., 2024), MATH-500 (Hendrycks et al., 2021a), Minerva (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). AMC (Li et al., 2024) comprises problems drawn from the AMC12 2022 and AMC12 2023 examinations. Moreover, when employing Qwen2.5-Math-7B as the backbone, we further conduct evaluations on GPQA-Diamond (Rein et al., 2024), a challenging and high-quality subset of the Graduate-Level Google-Proof Question Answering benchmark, as well as on ARC-c (Clark et al., 2018), an open-domain reasoning benchmark.
@@
where $P$ denotes model’s performance as introduced in Section 3.4 and $\gamma$ is the gate threshold. We fix $\gamma$ at 0 throughout all experiments on the Qwen Family models and $2$ for LLaMA, and provide relative ablation studies in Section 5.5. For hyperparameters, we use a constant learning rate of $5\times 10^{-6}$ and adopt the AdamW optimizer for the policy model. For rollout, we sample $8$ responses using a temperature of $1.0$. The maximum generation length is set to $8,192$ tokens for all other models. For other details that may not have been explicitly introduced, we have endeavored to follow previous works as closely as possible (Zhao et al., 2025; Zuo et al., 2025). All experiments were conducted on 8 x NVIDIA A800 80GB GPUs.
@@ 5.5 Gate Threshold Ablation
In this section, we investigate the effect of different gate thresholds $\gamma$. A value of $\gamma=0$ indicates that the model switches to SFT only when it fails all questions. Similarly, $\gamma=1$ and $\gamma=2$ correspond to settings where the model remains in on-policy reinforcement learning as long as it answers at least one or two out of eight questions correctly, respectively. To visualize the impact of the gating mechanism, we conduct experiments on the Qwen2.5-Math-1.5B under three different gate settings. As shown in Figure 8, we analyze the training dynamics by tracking the dynamics of rewards and the proportion of offline data utilized throughout training, thereby highlighting how different gate thresholds mediate the balance between leveraging offline demonstrations and incorporating online feedback. We observe that, under different gate thresholds, varying degrees of engagement with offline data–based SFT learning emerge. A larger gate threshold introduces a greater extent of SFT based on offline data, as expected.
@@
$\gamma=2$ $15.8$ $13.0$ $49.0$ $77.6$ $34.6$ $44.1$ $39.0$
$\gamma=1$ $\mathbf{18.1}$ $14.2$ $46.0$ $75.4$ $35.7$ $42.5$ $38.7$
@@
To further compare the performance across different gating strategies, we evaluate the three trained models on six benchmarks. Table 6 presents the results. Among the three configurations, $\gamma=0$ achieves the best overall performance with an average score of 41.9, outperforming both $\gamma=1$ (38.7) and $\gamma=2$ (39.0). This observation suggests that simply incorporating more SFT does not necessarily lead to better outcomes. Instead, it is crucial to maintain a dynamic balance between the exploration of RL and the exploitation of SFT. The optimal degree of this gating mechanism should be adjusted according to the characteristics of the base model and the specific training data employed.
@@ B.2 Extension: Adding a Trust-Region Regularizer
$\widehat{A}_{uni}^{(\lambda)}(\tau,q)=r(\tau\mid q)\;-\;\lambda\log\frac{\pi_{\theta}(\tau\mid q)}{\pi_{ref}(\tau\mid q)}\;+\;\mu\,\mathbb{1}\{\pi_{ref}=\pi_{\beta}\}\,\frac{\pi_{\beta}(\tau\mid q)}{\pi_{\theta}(\tau\mid q)}.$

### v2

@@ 1 Introduction
To circumvent these challenges, recent works have focused on integrating SFT or SFT-style imitation learning losses directly with RL objectives (Yan et al., 2025; Fu et al., 2025; Zhang et al., 2025). In these approaches, the model is updated using a composite loss function. The balance between the imitation and exploration components is governed by various strategies, including a fixed coefficient, a predefined schedule, a dynamic adjustment based on entropy, or a learnable parameter. These works predominantly treat the SFT and RL losses as two distinct objectives. And a detailed analysis of why these two learning signals can be effectively combined within a unified optimization process remains largely unexplored.
@@
Takeaways 1. UPGE provides a theoretical unification of a wide spectrum of post-training algorithms, covering both SFT and RL losses within a single formulation ($\S$ 2). 2. HPT is capable of outperforming previous post-training and mixed-policy algorithms across a diverse range of models ($\S$ 3). 3. Dynamic integration of SFT and RL in HPT achieves the highest Pass@1024, facilitating enhanced exploration and generalization of the model ($\S$ 4.1).
@@ 2 A Unified View on Post-Training Algorithms
GRPO (Shao et al., 2024) $\pi_{ref}=\pi_{\theta_{old}}$ $\hat{A}_{GRPO}=\frac{R(\tau_{j})-\text{mean}(\{R(\tau_{j})\}_{G_{on}})}{\text{std}(\{R(\tau_{j})\}_{G_{on}})}$ $\nabla\mathcal{J}_{GRPO}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{GRPO}\mathbb{1}_{\text{Clip}}}{\pi_{ref}(\tau)}$
REINFORCE (Ahmadian et al., 2024) $\pi_{ref}=\pi_{\theta}$ $\hat{A}_{REINFORCE}=\pm 1$ $\nabla\mathcal{J}_{REF.}(\theta)=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{REF.}}{\pi_{\theta}(\tau)}$
CISPO (Chen et al., 2025) $\pi_{ref}=\pi_{\theta_{old}}$ $\hat{A}_{CISPO}=\hat{A}_{GRPO}$ $\nabla\mathcal{J}_{CISPO}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{CISPO}\mathbb{1}_{\text{CIS-Mask}}}{\pi_{ref}(\tau)}$
GSPO (Zheng et al., 2025) $\pi_{ref}=\pi_{\theta}\left(\frac{\pi_{\theta_{old}}(\tau_{i,j}|q_{i})}{\pi_{\theta}(\tau_{i,j}|q_{i})}\right)^{1/|\tau_{i,j}|}$ $\hat{A}_{GSPO}=\hat{A}_{GRPO}$ $\nabla\mathcal{J}_{GSPO}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{GSPO}\mathbb{1}_{\text{Seq-Clip}}}{\pi_{ref}(\tau)}$
@@
SRFT (Offline) (Fu et al., 2025) $\pi_{ref}\equiv 1$ $\hat{A}_{SRFT}=\frac{R(\tau_{j})-\text{mean}(\{R(\tau_{j})\}_{G_{on}\cup G_{off}})}{\text{std}(\{R(\tau_{j})\}_{G_{on}\cup G_{off}})}$ $\nabla\mathcal{J}_{SRFT}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{SRFT}}{\pi_{ref}(\tau)=1}$
LUFFY (Offline) (Yan et al., 2025) $\pi_{ref}\equiv 1$ $\hat{A}_{LUFFY}=\hat{A}_{SRFT}$ $\nabla\mathcal{J}_{LUFFY}=\nabla\pi_{\theta}(\tau)\frac{\hat{A}_{LUFFY}}{\pi_{ref}(\tau)=1}f_{\text{shape}}^{\prime}$
@@ 2.2 Derivation of the Unified Policy Gradient Estimator
$\widehat{A}_{uni}(\tau,q)=\underbrace{r(\tau\mid q)}_{\widehat{A}_{\mathrm{RL}}(\tau,q)}\;+\;\underbrace{\mu\,\frac{\pi_{\beta}(\tau\mid q)}{\pi_{\theta}(\tau\mid q)}}_{\widehat{A}_{\mathrm{SFT}}(\tau,q)}.$ (4)
@@ 2.3 Gradient Component Analysis
Apart from the strong connection to data source and sampling polices, some studies employ a hand-crafted reweight factor within the reference policy denominator. These works (Yan et al., 2025; Zhang et al., 2025) typically find desirable token properties and purposefully place a higher/lower weight on these desirable/undesirable tokens, respectively.
@@
There are two commonly used settings for estimating the sequence-level advantage function: the fixed advantage setting and the adaptive advantage setting. The fixed setting considers $\hat{A}=\pm 1$ given the rule-based verification, which is adapted by REINFORCE and implicitly by SFT (where all sequences are positive samples). Alternatively, recent studies have focused on using adaptive advantage estimations, performing re-centering or normalization based on the performance of the current rollout group. Notably, GRPO and its variants, such as DAPO (Yu et al., 2025) and LUFFY (Yan et al., 2025), use unit normalization such that the advantage estimation of the group has a unit standard deviation. Other approaches, such as Dr. GRPO (Liu et al., 2025), RLOO (Ahmadian et al., 2024), and REINFORCE++ (Hu et al., 2025a), claim that dividing the standard deviation introduces a difficulty bias and that only recentering is adequate.
@@ 2.4 Hybrid Post-Training with Performance Feedback
 $P\leftarrow\frac{1}{n}\sum_{i=1}^{n}v(\tau_{i})$$\alpha\leftarrow f(P),\quad\beta\leftarrow g(P)$# Performance feedback on question $q$ Compute on-policy RL loss $\mathcal{L}_{\mathrm{RL}}$ using rollouts $\{\tau_{i}\}$ and normalized advantages derived from $\{R(\tau_{i})\}$. Compute SFT loss $\mathcal{L}_{\mathrm{SFT}}$ on the supervising trajectory $\tau^{\star}$. $\mathcal{L}\leftarrow\alpha\,\mathcal{L}_{\mathrm{RL}}+\beta\,\mathcal{L}_{\mathrm{SFT}}$# Mixed loss with performance feedback coefficients $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}$
@@ 3.1 Experimental Setup
We evaluate HPT on $6$ mathematical reasoning benchmarks: AIME 2024 (Li et al., 2024), AIME 2025 (Li et al., 2024), AMC (Li et al., 2024), MATH-500 (Hendrycks et al., 2021), Minerva (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). AMC (Li et al., 2024) comprises problems drawn from the AMC12 2022 and AMC12 2023 examinations. Moreover, when employing Qwen2.5-Math-7B as the backbone, we further conduct evaluations on GPQA-Diamond (Rein et al., 2024), a challenging and high-quality subset of the Graduate-Level Google-Proof Question Answering benchmark, as well as on ARC-c (Clark et al., 2018), an open-domain reasoning benchmark.
@@
where $P$ denotes model’s performance as introduced in Section 2.4 and $\gamma$ is the gate threshold. We fix $\gamma$ at 0 throughout all experiments on the Qwen Family models and $2/8$ for LLaMA, and provide relative ablation studies in Section 4.5. For hyperparameters, we use a constant learning rate of $5\times 10^{-6}$ and adopt the AdamW optimizer for the policy model. For rollout, we sample $8$ responses using a temperature of $1.0$. The maximum generation length is set to $8,192$ tokens for all other models. For other details that may not have been explicitly introduced, we have endeavored to follow previous works as closely as possible (Zhao et al., 2025; Zuo et al., 2025). All experiments were conducted on 8 x NVIDIA A800 80GB GPUs.
@@ 4.5 Gate Threshold Ablation
In this section, we investigate the effect of different gate thresholds $\gamma$. A value of $\gamma=0$ indicates that the model switches to SFT only when it fails all questions. Similarly, $\gamma=1/8$ and $\gamma=2/8$ correspond to settings where the model remains in on-policy reinforcement learning as long as it answers at least one or two out of eight questions correctly, respectively. To visualize the impact of the gating mechanism, we conduct experiments on the Qwen2.5-Math-1.5B under three different gate settings. As shown in Figure 8, we analyze the training dynamics by tracking the dynamics of rewards and the proportion of offline data utilized throughout training, thereby highlighting how different gate thresholds mediate the balance between leveraging offline demonstrations and incorporating online feedback. We observe that, under different gate thresholds, varying degrees of engagement with offline data–based SFT learning emerge. A larger gate threshold introduces a greater extent of SFT based on offline data, as expected.
@@
$\gamma=2/8$ $15.8$ $13.0$ $49.0$ $77.6$ $34.6$ $44.1$ $39.0$
$\gamma=1/8$ $\mathbf{18.1}$ $14.2$ $46.0$ $75.4$ $35.7$ $42.5$ $38.7$
@@
To further compare the performance across different gating strategies, we evaluate the three trained models on six benchmarks. Table 6 presents the results. Among the three configurations, $\gamma=0$ achieves the best overall performance with an average score of 41.9, outperforming both $\gamma=1/8$ (38.7) and $\gamma=2/8$ (39.0). This observation suggests that simply incorporating more SFT does not necessarily lead to better outcomes. Instead, it is crucial to maintain a dynamic balance between the exploration of RL and the exploitation of SFT. The optimal degree of this gating mechanism should be adjusted according to the characteristics of the base model and the specific training data employed.
@@ B.2 Extension: Adding a Trust-Region Regularizer
$\widehat{A}_{uni}^{(\lambda)}(\tau,q)=r(\tau\mid q)\;-\;\lambda\log\frac{\pi_{\theta}(\tau\mid q)}{\pi_{ref}(\tau\mid q)}\;+\;\mu\ \frac{\pi_{\beta}(\tau\mid q)}{\pi_{\theta}(\tau\mid q)}.$